# Reconstructing Materials Tetrahedron: Challenges in Materials Information Extraction

### v2

@@ 3.1.3 Extracting compositions from table and text jointly:
Extracting information from PI tables is more challenging than extracting from CI tables, as the incomplete information in the table regarding the composition should be inferred from the text. A detailed analysis of 50 PI tables revealed that 36% of the tables have unique challenges and are not ”regular”. To clarify this point further, we discuss some of these challenges below, while also defining a “regular” MCC-PI table in Fig 15.
@@
a. Unusual variables used: Other than the common variables like , , , , and , we also encounter variables like , , and in 4% of the manuscripts. Distinguishing some of them such as S or Y is difficult as they are valid symbols for chemical elements as well Henry and Hill (2003).
b. Variables representing composition in text not found in tables: A generic way of extracting the composition when an arithmetic equation of the composition containing variables is mentioned in the text is to connect it to the variables located in the headings of the table and substitute it with values mentioned under it. Extraction becomes difficult if there is an absence or mismatch between the variable name in the table and the text, as shown in Fig. 7(a). We found 8% of the tables posing this challenge Yoshida et al. (2003).
@@
c. Composition present across multiple columns: The composition of the material is spread across multiple columns/rows (instance depicted in Fig. 7(b) Kadono et al. (1999)), or the table does not follow any fixed orientation. This is observed among 4% of the PI tables.
@@
d. Composition partly in the table and partly in text: Although PI tables contain the composition partly, it is expected that the complete information is available in the text. But in rare occurrences, as depicted by Fig. 7(c), we observe that only the remaining part of the composition, that isn’t mentioned in the table, is present in the text. This makes linking the parts of compositions in the text and tables challenging. Thus, extracting the whole composition is extremely difficult; a case seen in less than 1% of the PI tables.
@@
e. Presence of multiple variables: 8% of the PI tables contains more than one variable, all of which need to be taken into account to extract the composition correctly. As discussed previously, variables can be of various forms, which might make the extraction of multiple variables a challenging task. Hayashi et al. (1998)
@@ 3.2 Extracting properties from tables
Until now, we focused on the extraction of compositions from tables. Now, we discuss the challenges with property extraction. To this extent, we analyzed 100 arbitrarily selected property tables. The observations based on this analysis are as follows.
@@
a. Semantically similar row/column headers: 19% of the tables have similar abbreviations or headers with similar descriptions for different properties. For example, in Fig. 8(a), the headings of the columns are T, T, T, T, T, T Gu et al. (2002). Identifying the desired property by a predictor model or someone without domain knowledge can be difficult in this case.
@@
b. Same property measured under different conditions: The same property can be measured with different techniques or under different conditions. Therefore, it is important to extract the correct contextual information related to the reported property. Some recurrent scenarios include witnessing tables with various refractive indexes (RIs) at different wavelengths Qiao et al. (2005) (see Fig. 11), glass transition temperatures at different heating rates Giridhar and Mahadevan (1998), or hardness at different testing loads. We encountered 9% of the property tables exhibiting this challenge.
@@
c. Information in caption/footer instead of tables: Often, properties are mentioned with abbreviations in the headings of tables, which are semantically close to other properties (for example, Fig. 8(a)). The information regarding their abbreviation is commonly found in the caption or footer of the table. We observed 30% of the tables having this characteristic Scudino et al. (2005); Deparis et al. (2005). Further, 2% of the tables have no information on the properties units. However, these are found in the caption and footer of the tables Okuno et al. (2005). Hence, text from these sections might be handy for extracting our desired properties.
@@
d. Property recorded under various acronyms: It is a common practice to record property names with their abbreviations. Some properties can have various abbreviations like density is represented with either or d, Young’s modulus with YM or E, and activation energy with E, Ae, or E.
@@
e. Identical acronyms representing different entities: We encountered tables (see Fig. 8(b,c)) where the commonly used acronyms are used to represent different entities; not the usual property they generally represent. For example, ‘n’, which is mostly used to represent RI, is also used to represent equation parameters specific to the experiments. Another commonly seen instance is ‘d’ which is used to represent density Seshadri et al. (2014) and has also been used to represent fractal bond connectivity Abd El-Moneim (2014), lattice parameters and equation parameters. This suggests that using a string-matching IE algorithm can result in poor performance in such cases.
@@
f. Range of values (min-max) given instead of mean value: In very few cases (< 1% tables), we encountered property values reported in range rather than a single value. For example, the values of T are reported in the range 930-945C Besson et al. (2000). Only a domain expert would know which value to take for a corresponding property between the min, max, or mean of the documented values. This might depend on the property or the application intended to be used, and will also be reflected in the IE algorithm.
3.3 Common for both composition and property extraction:
@@
a. Same composition or property represented with different units: Occasionally, tables are presented where the essential information is recorded in multiple units within the same table in different columns/rows. Fig. 9 shows a composition table having composition in both mol% and wt% Konishi et al. (2000), and a property table having glass transition temperature (T) mentioned in both C and K Rocherullé et al. (1998). This can lead to duplication of the extracted data.
b. Multiple ways of reporting the same unit: Despite the well-known and accepted conventions for writing the SI units Thompson and Taylor (2008), research publications resort to multiple ways of reporting the same unit. For instance, for g/cm, several variations are observed in peer-reviewed publications such as gm/cm, g.cm, g/cm, gcm, g/cc, gm/cc, gw/cm, gm cc. Similar observations are made for kg/m, where variations such as kgm, kg/m, kg m are presented. Extracting the correct unit and normalizing it to a standard form is an essential task. Thus, while there are standard rules for writing SI units, it is observed that these are not followed strictly in scientific publications.
3.4 IE or manufacturing and characterizing materials
@@
a. Precursor extraction: A research paper generally investigates materials of a similar kind. Hence, it has to be assumed that all the materials are manufactured using the same precursors. In research papers where batch composition is mentioned in tables, the challenges are similar as mentioned in Section 3.1.1(b). In papers where researchers discuss the patented materials, they refer to them by their trademark name, for example, Pyrex, BOROSIL, Gorilla, etc. and hence their precursor information is not provided. However, papers discussing materials reported in previous publications, provide references to those papers reporting the required information in detail.
@@
b. Processing conditions extraction: Processing conditions could be extremely non-linear and convoluted. Consider the set of sentences describing the processing conditions Roderick et al. (2001) as follows. “… powders were weighed and mixed thoroughly before being transferred to a 90 Pt/10 Rh crucible, heated at 320°C and maintained between 1000 and 1400°C depending on composition, for approximately 25 min. After annealing for approximately three hours, the glass was allowed to cool slowly to room temperature…”. Hence, the challenges here are to extract temperatures and duration for each process, like heating, annealing, and cooling, along with the environmental conditions and experimental apparatus. Sometimes, these conditions are also mentioned in the table (see Fig. 11), and their extraction poses similar challenges as described in Section 3.2(b).
@@
c. Testing conditions extraction: The testing conditions mainly comprise the sample characteristics, dimensions, test name, instrument name, instrument settings, and testing variables like temperature, wavelength, load, etc. Consider the following excerpt from Sigoli et al. (2001): “The porous microstructure of the matrix was investigated by scanning electron microscopy (SEM) (JEOL JSM T330A), by infrared spectroscopy (IR) in a FT-IR spectrometer (Perkin Elmer Spectrum 2000), and by X-ray powder diffractometry (XRD) (Siemens D-5000). The phase separation process was investigated by Raman microscope. The room temperature Raman measurements were performed through Raman imaging microscope (Renishaw) system 3000, with the 632.8 nm He–Ne laser line for excitation”. The boldface text indicates the information to be extracted for obtaining a complete understanding of the testing process of a material. Fig. 11 lists different wavelengths at which a material is tested to obtain wavelengths. The challenges faced in IE for this case will be similar to the ones posed in Section 3.2(b).
@@
Material structure: To study the structure of materials, researchers perform X-ray diffraction studies, obtain the Raman spectra, optical micrographs, and scanning electron micrographs depending upon the depth of detail about the material structure required. This information is mostly reported in figures and the figure description in the text provides some important details about the material’s structure. In the statement, “The Raman spectrum of the porous phase (Fig. 6(b)) shows only one band at 277 cm assigned to silica vibrations…”Sigoli et al. (2001), the information about Raman spectra is already shown in the graph, and the text mentions only critical findings.
@@
To summarise, the extraction of precursors, processing, and testing conditions from text poses challenges related to named entity recognition and relation extraction, which requires the need for specialized datasets and model development. Further, the extracted entities should be linked with the respective materials. The challenges faced during IE from tables for these variables require overcoming similar challenges as explained earlier for composition (Section 3.1.1) and properties (Section 3.2)
@@ 3.5 MatSci Knowledge-base: Linking extracted information
Material IDs are required to link information across multiple tables. For instance, in Fig. 12 Solvang et al. (2004), we obtain the composition of CAS1 from Table 1 and of this material from another table(Fig. 12 Table 2.). We detected 187 out of 2536 (7.37%) publications where inter-table IE is necessary and found difficulties in 81 of them while connecting the different components of the tetrahedron.
a. Different material IDs in different tables: The same materials have been reported with different IDs in different tables. 21 out of the 81 research papers (25.93%) have this challenge Romero et al. (2000); Martinelli et al. (2000).
b. Material IDs absent from tables: We detected a few MatSci papers where no IDs are present in the tables. There exist 23 out of 81 (28.40%) documents having this challenge Peitl and Zanotto (1999).
c. One of the tables does not contain material IDs: While connecting two tables, there are cases where IDs are mentioned only in one table Beggiora et al. (2003) (37 out of 81 (45.67%) papers with this challenge).
@@ 3.6 ID Analysis
As material ID is the key component in connecting materials from tables to text, across two different tables, or also across different sections of the text, we investigated arbitrarily selected 50 articles containing material IDs in the tables and recorded their semantic pattern to observe the semantics used by authors to refer to materials. We found that a majority of the authors prefer to use acronyms or self-made codes as IDs for referring to the materials, followed by natural numbers and standard material names, illustrated in Fig. 13. Material IDs are generally present at the beginning of the table and very rarely seen in the middle or end. Often, we come across tables having IDs which contain relevant information like the processing conditions of the material, or information about the state of the glass like amorphous or crystalline, or its composition, which are not separately mentioned in the table. These information are generally encoded as abbreviations, and extracting them can be challenging. In Table 1, we describe different cases along with the percentage occurrences. Note that the composition of the material should not be confused with its ID, as both are separate entities. An ID is expected to be shorter in length, most likely an acronym and unique to each material.
@@
| Article contains IDs Interconnected | 2 |
@@ 4 Guidelines for writing IE-friendly MatSci tables
Tables should be reported in such a way that automated extraction and the detection of the desired information are easy. Some of our suggestions are as follows (illustrated with Fig. 14, adapted from Kosuge et al. (1998)):
a. Use column orientation: Many IE algorithms which have been developed for tables have considered column orientation only. Also, we showed that 93% of the published tables are column-oriented. The following suggestions assume that we are following column orientation.
@@
c. Use proper and descriptive headers: The headers should contain the chemical formula of the compounds or elements which make up the materials, along with the acronyms of the reported properties, with processing, and testing conditions. If precursors, processing, and testing conditions are common, they can be omitted from tables.
@@
g. Consistent IDs: Material IDs should be formed as an acronym of its comprising constituents. They should be consistent in the whole article, i.e., there should not be more than one ID referring to the same material.
@@
j. Captions: All tables should have a clear, concise and descriptive caption. The table caption should clearly explain the acronyms used in the tables.
@@ 5 Conclusion and future work
The literature is replete with IE challenges and algorithms to extract information about materials. However, there exists no study which quantifies how much benefit can be obtained if a particular challenge is solved. In this paper, we have identified and quantified several unresolved challenges present in IE for every aspect of the MatSci tetrahedron. Specifically, we pointed out the locations in a MatSci research paper, where each piece of information on the MatSci tetrahedron of a given material is reported. Further, we outlined the challenges associated with IE and linking them to build the MatSci KB. We hope this extensive analysis will motivate researchers to focus on the challenges in the field, giving an insight into the gain associated with each of these challenges. This will also enable the researchers to identify the right problems to focus on based on the desired outcome. Finally, we provided recommendations towards an IE-friendly table format, which will enhance the automated extraction of the desired information and improve the tabular understanding of the researchers. Such concerted efforts are required to streamline the reporting in MatSci articles, thereby accelerating IE for materials discovery.

### v3

@@ 3.1.3 Extracting compositions from table and text jointly
Extracting information from PI tables is more challenging than extracting from CI tables, as the incomplete information in the table regarding the composition should be inferred from the text. A detailed analysis of 50 PI tables revealed that 36% of the tables have unique challenges and are not ”regular”. To clarify this point further, we discuss some of these challenges below while also defining a “regular” MCC-PI table in Fig 14. We have cross-checked all the reported challenges in this section by using the DiSCoMaT[8], the best IE model for composition extraction from MatSci tables [4], which also handles PI tables; and found that the model was unsuccessful in extracting composition from tables having these characteristics.
@@
a. Unusual variables used: Other than the common variables like , , , , and , we also encounter variables like , , and in 4% of the manuscripts. Distinguishing some of them, such as S or Y, is difficult as they are valid symbols for chemical elements as well [40].
@@
b. Composition present across multiple columns: The composition of the material is spread across multiple columns/rows (instance depicted in Fig. 6(a) [41]), or the table does not follow any fixed orientation. This is observed among 4% of the PI tables.
@@
c. Composition partly in the table and partly in text: Although PI tables contain the composition partly, it is expected that the complete information is available in the text. But in rare occurrences, as depicted by Fig. 6(b), we observe that only the remaining part of the composition, which is not mentioned in the table, is present in the text. This makes linking the parts of compositions in the text and tables challenging. Thus, extracting the whole composition is extremely difficult, a case seen in less than 1% of the PI tables [42].
@@
d. Presence of multiple variables: We found 6% of the PI tables having more than one variable, all of which need to be taken into account to extract the composition correctly. As discussed previously, variables can be of various forms, making extracting multiple variables a challenging task [43, 44].
@@ 3.2 Extracting properties from tables
Until now, we focused on the extraction of compositions from tables and text. In this section, we discuss the challenges with property extraction. To this extent, we analyzed 100 arbitrarily selected property tables. The observations based on this analysis are as follows.
@@
a. Semantically similar row/column headers: 19% of the tables have similar abbreviations or headers with similar descriptions for different properties. For example, in Fig. 7(a), the headings of the columns are Tg, Tx1, Tx2, Tx3, Tx, Tm [45]. Identifying the desired property by a predictor model or someone without domain knowledge can be difficult in this case.
@@
b. The same property measured under different conditions: The same property can be measured with different techniques or under different conditions. Therefore, it is important to extract the correct contextual information related to the reported property. Some recurrent scenarios include witnessing tables with various refractive indexes (RIs) at different wavelengths [46] (see Fig. 10), glass transition temperatures at different heating rates [47], or hardness at different testing loads. We encountered 9% of the property tables exhibiting this challenge.
@@
c. Information in caption/footer instead of tables: Often, properties are mentioned with abbreviations in the headings of tables, which are semantically close to other properties (for example, Fig. 7(a)). The information regarding their abbreviation is commonly found in the caption or footer of the table. We observed 30% of the tables having this characteristic [48, 49]. Further, 2% of the tables have no information on the properties units. However, these are found in the caption or footer of the tables [50]. Hence, text from these sections might be handy for extracting our desired properties.
@@
d. Property recorded under various acronyms: It is a common practice to record property names with their abbreviations. Some properties can have various abbreviations like density is represented with either or d, Young’s modulus with YM or E, and activation energy with E0, Ae, or Ea.
@@
e. Identical acronyms representing different entities: We encountered tables (see Fig. 7(b,c)) where the commonly used acronyms are used to represent different entities; not the usual property they generally represent. For example, ‘n’, which is mostly used to represent RI, is also used to represent equation parameters specific to the experiments. Another commonly seen instance is ‘d’ which is used to represent density [51] and has also been used to represent fractal bond connectivity [52], lattice parameters, and equation parameters. This suggests that using a string-matching IE algorithm can result in poor performance in such cases.
@@
f. Range of values (min-max) given instead of mean value: In very few cases (< 1% tables), we encountered property values reported in range rather than a single value. For example, the values of Tg are reported in the range 930-945∘C [53]. Only a domain expert would know which value to take for a corresponding property between the min, max, or mean of the documented values. This might depend on the property or the application intended to be used, and will also be reflected in the IE algorithm.
3.3 Challenges common for both composition and property extraction:
@@
a. Same composition or property represented with different units: Tables are sometimes (2%) presented with the essential information recorded in multiple units in different columns/rows. Fig. 8 shows a composition table having composition in both mol% and wt% [54], and a property table having glass transition temperature (Tg) mentioned in both ∘C and K [55]. This can lead to duplication of the extracted data.
b. Multiple ways of reporting the same unit: Despite the well-known and accepted conventions for writing the SI units [56], research publications resort to multiple ways of reporting the same unit. For instance, for g/cm3, several variations are observed in peer-reviewed publications such as gm/cm3, g.cm-3, g/cm3, gcm-3, g/cc, gm/cc, gw/cm3, gm cc-1. Similar observations are made for kg/m3, where variations such as kgm-3, kg/m3, kg m-3 are presented. Extracting the correct unit and normalizing it to a standard form is an essential task. Thus, while there are standard rules for writing SI units, it is observed that these are not strictly followed in scientific publications.
c. Multiple tables merged in one: A rarely seen challenge (<1%) is illustrated in Fig. 9, where many tables are concatenated in a long or broad table, which leads to difficulties in extracting the required details [57].
Note that none of these challenges could be solved using the IE model DiSCoMaT [8] and GPT-4.
3.4 IE for manufacturing and characterizing materials
@@
a. Precursor extraction: A research paper generally investigates materials of a similar kind. Hence, it has to be assumed that all the materials are manufactured using the same precursors. In research papers where batch composition is mentioned in tables, the challenges are similar as mentioned in Section 3.1.1(b). In papers where researchers discuss the patented materials, they refer to them by their trademark name, for example, Pyrex, BOROSIL, Gorilla, etc., and hence their precursor information is not provided. However, papers discussing materials reported in previous publications, provide references to those papers reporting the required information in detail.
@@
b. Processing conditions extraction: Processing conditions reporting could be extremely non-linear and convoluted. Consider the set of sentences describing the processing conditions [58] as follows. “… powders were weighed and mixed thoroughly before being transferred to a 90 Pt/10 Rh crucible, heated at 320°C and maintained between 1000 and 1400°C depending on composition, for approximately 25 min. After annealing for approximately three hours, the glass was allowed to cool slowly to room temperature…”. Hence, the challenges here are to extract temperatures and duration for each process, like heating, annealing, and cooling, along with the environmental conditions and experimental apparatus. Sometimes, these conditions are also mentioned in the table (see Fig. 10), and their extraction poses similar challenges as described in Section 3.2(b).
@@
c. Testing conditions extraction: The testing conditions mainly comprise the sample characteristics, dimensions, test name, instrument name, instrument settings, and testing variables like temperature, wavelength, load, frequency, pressure, etc. Consider the following excerpt from [59]: “The porous microstructure of the matrix was investigated by scanning electron microscopy (SEM) (JEOL JSM T330A), by infrared spectroscopy (IR) in a FT-IR spectrometer (Perkin Elmer Spectrum 2000), and by X-ray powder diffractometry (XRD) (Siemens D-5000). The phase separation process was investigated by Raman microscope. The room temperature Raman measurements were performed through Raman imaging microscope (Renishaw) system 3000, with the 632.8 nm He–Ne laser line for excitation”. The boldface text indicates the information to be extracted for obtaining a complete understanding of the testing process of a material. Fig. 10 lists different wavelengths at which a material is tested to obtain refractive index. The challenges faced in IE for this case will be similar to the ones posed in Section 3.2(b).
@@
Material structure: To study the structure of materials, researchers perform X-ray diffraction studies, obtain the Raman spectra, optical micrographs, and scanning electron micrographs depending upon the depth of detail about the material structure required. This information is mostly reported in figures and the figure description in the text provides some important details about the material’s structure. In the statement, “The Raman spectrum of the porous phase (Fig. 6(b)) shows only one band at 277 cm-1 assigned to silica vibrations…”[59], the information about Raman spectra is already shown in the graph, and the text mentions only critical findings.
@@
To summarise, the extraction of precursors, processing, and testing conditions from text poses challenges related to named entity recognition and relation extraction, which requires the need for specialized datasets and model development. there exist several materials science domain-specific models capable of extracting this information but their performance (F1-Score) on different types of desired entities ranges from as low as 33% [5] (interlayer materials for batteries, taken from SOFC dataset[60]) to 93% [61](materials tag, taken from MatScholar dataset[62]). There also exist some knowledge graphs created using these tools like MatKG[63], however, the quality of the information in such sources is as good as the underlying model. Further, on relation-extraction tasks, the best-performing models have an F1 score of 0.82[64], which indicates significant efforts required to facilitate the information extraction and complete materials science tetrahedron. Further, the extracted entities should be linked with the respective materials. The challenges faced during IE from tables for processing and testing variables require overcoming similar challenges as explained earlier for composition (Section 3.1.1) and properties (Section 3.2).
@@ 3.5 MatSci Knowledge-base: Linking extracted information
Material IDs are required to link information across multiple tables. For instance, in Fig. 11 [65], we obtain the composition of CAS1 from Table 1 and of this material from another table(Fig. 11 Table 2.). Every material in an article should have a unique ID, which should be used consistently across the whole article to denote the corresponding material. Any exception to this will lead to difficulties in linking our extracted information. We detected 187 out of 2536 (7.37%) publications where inter-table IE is necessary and found difficulties in 81 of them while connecting the different components of the tetrahedron.
a. Different material IDs in different tables: The same materials have been reported with different IDs in different tables. 21 out of the 81 research papers (25.93%) have this challenge [66, 67].
b. Material IDs absent from tables: We detected MatSci papers where no IDs are present in the tables. There exist 23 out of 81 (28.40%) documents having this challenge, where compositions of the materials and their corresponding properties are reported in separate tables, but neither of the tables have any ID present denoting the material. [68].
c. One of the tables does not contain material IDs: While connecting two tables, there are cases where IDs are mentioned only in one table [69] (37 out of 81 (45.67%) papers with this challenge).
@@ 3.6 ID Analysis
As material ID is the key component in connecting materials from tables to text, across two different tables, or also across different sections of the text, we investigated arbitrarily selected 50 articles containing material IDs in the tables and recorded their semantic pattern to observe the semantics used by authors to refer to materials. We found that a majority of the authors prefer to use acronyms or self-made codes as IDs for referring to the materials, followed by natural numbers and standard material names, illustrated in Fig. 12. Material IDs are generally present at the beginning of the table and very rarely seen in the middle or end. Often, we come across tables having IDs that contain relevant information like the processing conditions of the material, or information about the state of the glass like amorphous or crystalline, or its composition, which are not separately mentioned in the table. These information are generally encoded as abbreviations, and extracting them can be challenging. In Table 1, we describe different cases along with the percentage occurrences. Note that the composition of the material should not be confused with its ID, as both are separate entities. An ID is expected to be shorter in length, most likely an acronym, and unique to each material.
@@
| The article contains IDs interconnected | 2 |
@@ 4 Guidelines for writing IE-friendly MatSci tables
Tables should be reported in such a way that automated extraction and the detection of the desired information are easy. Some of our suggestions are as follows (illustrated with Fig. 13, adapted from [70]):
a. Use column orientation: Many IE algorithms that have been developed for tables have considered column orientation only. Also, we showed that 93% of the published tables are column-oriented. The following suggestions assume that we are following column orientation.
@@
c. Use proper and descriptive headers: The headers should contain the chemical formula of the compounds or elements that make up the materials, along with the acronyms of the reported properties, with processing, and testing conditions. If precursors, processing, and testing conditions are common, they can be omitted from tables.
@@
g. Consistent IDs: Material IDs should be formed as an acronym of its comprising constituents. They should be consistent in the whole article, that is, there should not be more than one ID referring to the same material.
@@
j. Captions: All tables should have a clear, concise, and descriptive caption. The table caption should clearly explain the acronyms used in the tables.
@@ 5 Conclusion and future work
The literature is replete with IE challenges and algorithms to extract information about materials. However, there exists no study that quantifies how much benefit can be obtained if a particular challenge is solved. In this paper, we have identified and quantified several unresolved challenges present in IE for every aspect of the MatSci tetrahedron. Specifically, we pointed out the locations in a MatSci research paper where each piece of information on the MatSci tetrahedron of a given material is reported. Further, we outlined the challenges associated with IE and linking them to build the MatSci KB. We hope this extensive analysis will motivate researchers to focus on the challenges in the field, giving an insight into the gain associated with each of these challenges. This will also enable the researchers to identify the right problems to focus on based on the desired outcome. Finally, we provided recommendations for an IE-friendly table format to enhance the automated extraction of the desired information and improve the researchers’ tabular understanding. Such concerted efforts are required to streamline the reporting in MatSci articles, thereby accelerating IE for materials discovery.